# Can SSD-Mamba2 Unlock Reinforcement Learning for End-to-End Motion Control?

### v1

@@ Abstract
End-to-end reinforcement learning for motion control promises unified perception–action policies that scale across embodiments and tasks, yet most deployed controllers are either blind (proprioception-only) or rely on fusion backbones with unfavorable compute–memory trade-offs. Recurrent controllers struggle with long-horizon credit assignment, and Transformer-based fusion incurs quadratic cost in token length, limiting temporal and spatial context. We present a vision-driven cross-modal RL framework built on SSD-Mamba2, a selective state-space backbone that applies state-space duality (SSD) to enable both recurrent and convolutional scanning with hardware-aware streaming and near-linear scaling. Proprioceptive states and exteroceptive observations (e.g., depth tokens) are encoded into compact tokens and fused by stacked SSD-Mamba2 layers. The selective state-space updates retain long-range dependencies with markedly lower latency and memory use than quadratic self-attention, enabling longer look-ahead, higher token resolution, and stable training under limited compute. Policies are trained end-to-end under curricula that randomize terrain and appearance and progressively increase scene complexity. A compact, state-centric reward balances task progress, energy efficiency, and safety. Across diverse motion-control scenarios, our approach consistently surpasses strong state-of-the-art baselines in return, safety (collisions and falls), and sample efficiency, while converging faster at the same compute budget. These results suggest that SSD-Mamba2 provides a practical fusion backbone for scalable, foresightful, and efficient end-to-end motion control.
Index Terms:
Artificial Intelligence, Machine Vision, Learning Systems, Road Vehicles, Cognitive Science
I Introduction
@@
Quadrupedal robots are increasingly deployed in domains where safe and reliable mobility is critical, including inspection, disaster response, and exploration of unstructured terrains [1]. Unlike wheeled platforms, they can traverse rubble, stairs, and deformable substrates, but their control remains a longstanding challenge: locomotion must be robust under diverse dynamics, foresightful to anticipate hazards, and computationally efficient for deployment on embedded hardware [2].
@@
Deep reinforcement learning (DRL) offers a promising route toward such capabilities by training policies end-to-end from interaction. While DRL has enabled agile behaviors across terrains and disturbances, most prior controllers fall short in one of two ways. Proprioception-only (“blind”) agents can maintain stability but lack foresight, reacting only after contact with obstacles [3]. In contrast, cross-modal approaches that fuse vision with proprioception provide anticipatory information but rely on backbones with unfavorable compute–memory trade-offs [4]. Recurrent models struggle to capture long-horizon dependencies [5], hierarchical schemes complicate optimization [6], and Transformers incur quadratic costs in token length, limiting their ability to scale [7].
These limitations raise a broader question: Can recent advances in sequence modeling unlock practical end-to-end RL for motion control? To answer this, we explore the State Space Duality (SSD) formulation of Mamba2 [8], a selective state-space architecture that achieves near-linear time and memory complexity while retaining the ability to model long-range dependencies. Unlike attention-based fusion, SSD-Mamba2 performs hardware-aware recurrent scanning, providing scalability without sacrificing foresight. This property makes it a compelling candidate for quadrupedal locomotion, where policies must integrate both proprioceptive states and depth observations in real time.
@@
In this paper, we introduce a vision-driven cross-modal RL framework built on SSD-Mamba2. Proprioceptive states are embedded by a lightweight multilayer perceptron (MLP) [9], depth images are patchified into spatial tokens by a compact convolutional neural network (CNN) [10], and the resulting token sequence is fused by stacked SSD-Mamba2 layers. Policies and value functions are then trained end-to-end with Proximal Policy Optimization (PPO) [11] under domain randomization and an obstacle-density curriculum. A compact state-centric reward further encourages progress, energy efficiency, and safety [12].
@@
The remainder of this paper is organized as follows. Section II reviews related work on blind and cross-modal locomotion. Section III presents the methodology, including the problem formulation, the SSD-Mamba2 fusion backbone, and PPO optimization. Section IV describes the experimental setup and reports results, ablations, and analyses of efficiency, stability, and generalization. Section V concludes the paper and discusses future research directions.
II Related Work
II-A Quadrupedal Locomotion Control
@@
Research on quadrupedal locomotion has traditionally followed two paradigms: model-based control and learning-based methods. Early rule-based approaches relied on template dynamics, heuristics, and hand-engineered gait patterns for balance and foot placement [13, 14, 15, 16], providing insights into leg coordination but limited adaptability to diverse environments. Later, model-based pipelines advanced to constrained optimal control, including trajectory optimization over centroidal or full-body dynamics and model predictive control (MPC) with explicit contact constraints and task hierarchies [17, 18, 19, 20]. These methods achieved robust disturbance rejection and stable locomotion in structured settings, with examples such as MPC-based whole-body controllers for stair climbing and rapid trotting [21, 22, 23]. However, such controllers require accurate models of dynamics, ground contact, and friction, along with significant manual cost tuning. This reliance limits their scalability to unstructured or rapidly changing terrains, such as rubble or deformable substrates.
@@
By contrast, learning-based methods aim to reduce modeling effort by training adaptive policies directly from interaction. Model-free deep reinforcement learning (DRL) in particular has shown strong potential for quadrupeds in diverse scenarios [24, 25, 26]. With large-scale domain randomization, policies trained in simulation have been successfully transferred to hardware [27, 28]. Recent work has further improved adaptability, such as rapid motor adaptation (RMA) [29], which augments a base policy with an adaptation module to handle payload changes and novel terrains online. DRL policies have also demonstrated robustness under water, mud, and snow conditions [30, 31]. Despite these advances, the majority of policies remain proprioception-only. Blind controllers can recover from disturbances but lack foresight and react only after contact events [3], which limits proactive obstacle avoidance and foothold planning. This motivates incorporating exteroceptive sensing such as vision to endow quadrupeds with anticipatory capability.
II-B Vision-Driven Reinforcement Learning
@@
Vision has emerged as a powerful complement to proprioception by providing spatial foresight. Visual information enables robots to perceive obstacles before contact, anticipate irregular terrain, and plan footholds or body trajectories proactively. Several studies have incorporated visual input directly into DRL pipelines. For instance, [yu2021visual] showed that depth-augmented RL policies could traverse rough terrain with greater robustness than proprioceptive baselines. [duan2024learning] combined heightmap observations with a depth-to-heightmap predictor, enabling sim-to-real transfer on hardware. ViTAL [32] used vision for terrain-aware foothold selection and pose adaptation, improving safety on irregular surfaces. Other works explored semantic segmentation or heightmap predictors to enhance terrain awareness [4]. Hierarchical designs [6] further decomposed vision-based locomotion into high-level planning and low-level control.
@@
However, fusion strategies for vision and proprioception present trade-offs. Simple feature concatenation underuses spatial structure and temporal dependencies, leading to limited generalization. Hierarchical formulations complicate optimization and can propagate errors across levels. Transformer-based fusion improves foresight but incurs quadratic time and memory scaling in token length, limiting input resolution and training efficiency [7, 33]. As a result, despite their promise, many vision-driven RL frameworks remain difficult to scale or deploy efficiently. This motivates the search for backbones that retain foresight while being more computationally tractable.
II-C Sequence Models for Cross-Modal Fusion
@@
A central challenge in vision-driven RL is the design of effective cross-modal fusion backbones. Recurrent neural network (RNN)–based approaches, such as long short-term memory (LSTM) and gated recurrent unit (GRU) architectures, have been used to integrate visual and proprioceptive information for quadrupedal locomotion [34, 35]. These methods capture temporal context but suffer from vanishing gradients and limited long-horizon memory, especially in extended locomotion sequences. Transformer-based models have recently become popular for cross-modal fusion [7], offering strong representational power and long-range context modeling. However, Transformers scale quadratically with sequence length, which constrains spatial resolution, temporal context, and training efficiency. This makes them costly for real-time quadrupedal control, where compute and memory budgets are limited.
@@
State-space models (SSMs) offer an alternative by updating compact recurrent states with near-linear complexity [36]. The S4 model [37] first demonstrated efficient long-range dependency modeling on long-sequence benchmarks. Building on this, Mamba [38] introduced selective state updates with hardware-aware parallel scanning, achieving linear-time sequence modeling with strong throughput and memory efficiency. Variants such as AlignMamba [39], FusionMamba [40], and DepMamba [41] have extended this paradigm to multimodal tasks, showing improvements in alignment, fusion, and cross-modal consistency. More recently, Mamba2 advanced this line of work through the SSD formulation, which further enhances stability and expressivity while retaining near-linear scaling [8]. SSD-Mamba2 combines the efficiency of selective scanning with dual state representations, enabling more reliable modeling of long-horizon dependencies.
@@ III-A Overview
Figure 1: Overall architecture of the proposed framework. Proprioception and depth are encoded into tokens by MLP/CNN, fused by a SSD-Mamba2 backbone, and optimized with PPO.
III-B Problem Formulation
1) Markov Decision Process: The motion-control task is formulated as a Markov Decision Process (MDP) [33] $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{P}(s_{t+1}\mid s_{t},a_{t})$ is the transition distribution defined by the physics simulator, and $r(s_{t},a_{t})$ is the reward function. The goal is to learn a policy $\pi_{\theta}(a_{t}\mid s_{t})$ that maximizes the expected discounted return:
@@ III-B Problem Formulation
3) Action space: The action $a_{t}\in\mathbb{R}^{12}$ specifies the desired joint position targets for the robot’s 12 actuated joints. A proportional–derivative (PD) controller [42] converts these position targets into torque commands applied to the actuators.
@@
III-C SSD-Mamba2 Fusion Backbone
@@
Figure 2: Schematic of a Mamba2 SSM Layer.
@@
III-D Policy Optimization with PPO
PPO is a widely used model-free reinforcement learning algorithm for continuous control due to its stability and simplicity. Instead of performing large, potentially destabilizing policy updates, PPO restricts the update step with a clipping mechanism, ensuring stable and efficient training [11]. The objective of PPO is defined as:
@@ III-D Policy Optimization with PPO
IV Experiments
IV-A Environment Setup
All experiments are carried out on a laptop with an Intel Xeon(R) Platinum 8358P CPU (128 cores, 2.6 GHz base frequency) and a NVIDIA GeForce RTX 3090 GPU (24 GB, CUDA 12.8). The computing environment is based on Ubuntu 22.04, with physics simulation executed in PyBullet [43]. All models are developed in Python 3.8 using PyTorch 2.4.1.
@@ IV-A Environment Setup
(a) Thin Obstacles with Goals
(b) Thin Obstacle with Rugged Terrain
(c) Sphere Obstacles with Goals.
Figure 3: Simulated environments. Obstacle layouts are randomized at reset; all obstacle positions are static during an episode.
IV-B Baseline Methods

### v2

# 1]organization=School of Computer Science and Statistics, Trinity College Dublin, city=Dublin, country=Ireland

@@ Abstract
End-to-end reinforcement learning (RL) for motion control trains policies directly from sensor inputs to motor commands, enabling unified controllers for different robots and tasks. However, most existing methods are either blind (proprioception-only) or rely on fusion backbones with unfavorable compute–memory trade-offs. Recurrent controllers struggle with long-horizon credit assignment, and Transformer-based fusion incurs quadratic cost in token length, limiting temporal and spatial context. We present a vision-driven cross-modal RL framework built on SSD-Mamba2, a selective state-space backbone that applies state-space duality (SSD) to enable both recurrent and convolutional scanning with hardware-aware streaming and near-linear scaling. Proprioceptive states and exteroceptive observations (e.g., depth tokens) are encoded into compact tokens and fused by stacked SSD-Mamba2 layers. The selective state-space updates retain long-range dependencies with markedly lower latency and memory use than quadratic self-attention, enabling longer look-ahead, higher token resolution, and stable training under limited compute. Policies are trained end-to-end under curricula that randomize terrain and appearance and progressively increase scene complexity. A compact, state-centric reward balances task progress, energy efficiency, and safety. Across diverse motion-control scenarios, our approach consistently surpasses strong state-of-the-art baselines in return, safety (collisions and falls), and sample efficiency, while converging faster at the same compute budget. These results suggest that SSD-Mamba2 provides a practical fusion backbone for resource-constrained robotic and autonomous systems in engineering informatics applications.
keywords:
[…]
{highlights}
Establishing SSD-Mamba2 as a backbone for RL motion control. We present the first study to leverage the SSD–based Mamba2 architecture in reinforcement learning for end-to-end motion control, demonstrating its suitability as an efficient and scalable sequence model in this domain.
1 Introduction
@@
Quadrupedal robots are increasingly deployed in domains where safe and reliable mobility is critical, including inspection, disaster response, and exploration of unstructured terrains [imambi2021pytorch]. Unlike wheeled platforms, they can traverse rubble, stairs, and deformable substrates, but their control remains a longstanding challenge: locomotion must be robust under diverse dynamics, foresightful to anticipate hazards, and computationally efficient for deployment on embedded hardware [carpentier2021recent].
@@
Deep reinforcement learning (DRL) offers a promising route toward such capabilities by training policies end-to-end from interaction. While DRL has enabled agile behaviors across terrains and disturbances, most prior controllers fall short in one of two ways. Proprioception-only (“blind”) agents can maintain stability but lack foresight, reacting only after contact with obstacles [xie2022glide]. In contrast, cross-modal approaches that fuse vision with proprioception provide anticipatory information but rely on backbones with unfavorable compute–memory trade-offs [han2025multimodal]. Recurrent models struggle to capture long-horizon dependencies [li2024rapid], hierarchical schemes complicate optimization [jain2019hierarchical], and Transformers incur quadratic costs in token length, limiting their ability to scale [yang2021learning]. These limitations raise a broader question: Can recent advances in sequence modeling unlock practical end-to-end RL for motion control?
Recent state-space models (SSMs) provide an appealing alternative by combining recurrent structure with efficient sequence processing, but different SSM variants exhibit distinct trade-offs when applied to RL. Structured long-convolution models such as S4 [gu2021efficiently] rely on explicit continuous-to-discrete parameterizations and long convolution kernels, often involving additional design choices such as kernel truncation and fast Fourier transform–based implementations [beaudoin2002accurate]. While effective for offline or supervised sequence modeling, these design elements can increase hyperparameter sensitivity and tuning complexity in RL settings with long rollouts and non-stationary data distributions. Selective SSMs such as Mamba [gu2023mamba] address some of these challenges through input-dependent state updates and efficient scan operations. However, in Mamba, the link between recurrent state updates and their scan-based implementation remains implicit. In contrast, SSD-Mamba2 [dao2024transformersssmsgeneralizedmodels] is built on an explicit state-space duality (SSD) formulation, in which recurrent evolution and block-wise parallel scanning are derived from the same state-space representation. This explicit mapping enables a unified and implementation-efficient realization of recurrent computation, reducing design complexity and improving practical robustness [ibrahim2025survey].
These properties make SSD-Mamba2 well-suited for real-time robotic motion control. Its hardware-aware recurrent scanning achieves near-linear time and memory complexity while supporting streaming execution and long-range temporal modeling. This balance between recurrent stability and computational efficiency is critical for quadrupedal locomotion, where policies must continuously fuse proprioceptive feedback and depth observations under strict real-time and resource constraints.
@@
In this paper, we introduce a vision-driven cross-modal RL framework built on SSD-Mamba2. Proprioceptive states are embedded by a lightweight multilayer perceptron (MLP) [tolstikhin2021mlpmixerallmlparchitecturevision], depth images are patchified into spatial tokens by a compact convolutional neural network (CNN) [oshea2015introductionconvolutionalneuralnetworks], and the resulting token sequence is fused by stacked SSD-Mamba2 layers. Policies and value functions are then trained end-to-end with Proximal Policy Optimization (PPO) [schulman2017proximal] under domain randomization and an obstacle-density curriculum. A compact state-centric reward further encourages progress, energy efficiency, and safety [yang2022learning].
@@
The remainder of this paper is organized as follows. Section 2 reviews related work on blind and cross-modal locomotion. Section 3 presents the methodology, including the problem formulation, the SSD-Mamba2 fusion backbone, and PPO optimization. Section 4 describes the experimental setup and implementation details. Section 5 presents the experimental results and ablation studies. Section 6 concludes the paper and discusses directions for future research.
2 Related Work
2.1 Quadrupedal Locomotion Control
@@
Research on quadrupedal locomotion has traditionally followed two paradigms: model-based control and learning-based methods. Early rule-based approaches relied on template dynamics, heuristics, and hand-engineered gait patterns for balance and foot placement [miura1984dynamic, liu2014planning, habu2018simple, bledt2018cheetah], providing insights into leg coordination but limited adaptability to diverse environments. Later, model-based pipelines advanced to constrained optimal control, including trajectory optimization over centroidal or full-body dynamics and model predictive control (MPC) with explicit contact constraints and task hierarchies [grandia2019feedback, yao2021hierarchy, amatucci2024accelerating, elobaid2025adaptive]. These methods achieved robust disturbance rejection and stable locomotion in structured settings, with examples such as MPC-based whole-body controllers for stair climbing and rapid trotting [di2018dynamic, ding2019real, carius2019trajectory]. However, such controllers require accurate models of dynamics, ground contact, and friction, along with significant manual cost tuning. This reliance limits their scalability to unstructured or rapidly changing terrains, such as rubble or deformable substrates.
@@
By contrast, learning-based methods aim to reduce modeling effort by training adaptive policies directly from interaction [wang2026locomamba]. Model-free DRL in particular has shown strong potential for quadrupeds in diverse scenarios [li2021reinforcement, zhang2021terrain, bussola2025guided]. With large-scale domain randomization, policies trained in simulation have been successfully transferred to hardware [tan2018sim, hwangbo2019learning]. Recent work has further improved adaptability, such as rapid motor adaptation [kumar2021rma, wang2025quadkan], which augments a base policy with an adaptation module to handle payload changes and novel terrains online. DRL policies have also demonstrated robustness under water, mud, and snow conditions [lee2020learning, xie2021dynamics, wang2025humam]. Despite these advances, the majority of policies remain proprioception-only. Blind controllers can recover from disturbances but lack foresight and react only after contact events [xie2022glide], which limits proactive obstacle avoidance and foothold planning. This motivates incorporating exteroceptive sensing such as vision to endow quadrupeds with anticipatory capability.
2.2 Vision-Driven Reinforcement Learning
@@
Vision has emerged as a powerful complement to proprioception by providing spatial foresight. Visual information enables robots to perceive obstacles before contact, anticipate irregular terrain, and plan footholds or body trajectories proactively. Several studies have incorporated visual input directly into DRL pipelines. For instance, yu2021visual showed that depth-augmented RL policies could traverse rough terrain with greater robustness than proprioceptive baselines. duan2024learning combined heightmap observations with a depth-to-heightmap predictor, enabling sim-to-real transfer on hardware. ViTAL [fahmi2022vital] used vision for terrain-aware foothold selection and pose adaptation, improving safety on irregular surfaces. Other works explored semantic segmentation or heightmap predictors to enhance terrain awareness [han2025multimodal]. Hierarchical designs [jain2019hierarchical] further decomposed vision-based locomotion into high-level planning and low-level control.
@@
However, fusion strategies for vision and proprioception present trade-offs. Simple feature concatenation underuses spatial structure and temporal dependencies, leading to limited generalization. Hierarchical formulations complicate optimization and can propagate errors across levels. Transformer-based fusion improves foresight but incurs quadratic time and memory scaling in token length, limiting input resolution and training efficiency [yang2021learning, singh2022reinforcement]. As a result, despite their promise, many vision-driven RL frameworks remain difficult to scale or deploy efficiently. This motivates the search for backbones that retain foresight while being more computationally tractable.
2.3 Sequence Models for Cross-Modal Fusion
@@
A central challenge in vision-driven RL is the design of effective cross-modal fusion backbones. Recurrent neural network (RNN)–based approaches, such as long short-term memory (LSTM) and gated recurrent unit (GRU) architectures, have been used to integrate visual and proprioceptive information for quadrupedal locomotion [xiao2024egocentric, lai2024world]. These methods capture temporal context but suffer from vanishing gradients and limited long-horizon memory, especially in extended locomotion sequences. Transformer-based models have recently become popular for cross-modal fusion [yang2021learning], offering strong representational power and long-range context modeling. However, Transformers scale quadratically with sequence length, which constrains spatial resolution, temporal context, and training efficiency. This makes them costly for real-time quadrupedal control, where compute and memory budgets are limited.
@@
State-space models (SSMs) offer an alternative by updating compact recurrent states with near-linear complexity [wang2026locomamba]. The S4 model [gu2021efficiently] first demonstrated efficient long-range dependency modeling on long-sequence benchmarks. Building on this, Mamba [gu2023mamba] introduced selective state updates with hardware-aware parallel scanning, achieving linear-time sequence modeling with strong throughput and memory efficiency. Variants such as AlignMamba [li2025alignmamba], FusionMamba [xie2024fusionmamba], and DepMamba [ye2025depmamba] have extended this paradigm to multimodal tasks, showing improvements in alignment, fusion, and cross-modal consistency. More recently, Mamba2 advanced this line of work through the SSD formulation, which further enhances stability and expressivity while retaining near-linear scaling [dao2024transformersssmsgeneralizedmodels]. SSD-Mamba2 combines the efficiency of selective scanning with dual state representations, enabling more reliable modeling of long-horizon dependencies.
@@ 3.1 Overview
Figure 1: Overview of the proposed architecture. Proprioception and depth are encoded into tokens by MLP/CNN, fused by a SSD-Mamba2 backbone, and optimized with PPO.
3.2 Problem Formulation
1) Markov Decision Process: The motion-control task is formulated as a Markov Decision Process (MDP) [singh2022reinforcement] $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{P}(s_{t+1}\mid s_{t},a_{t})$ is the transition distribution defined by the physics simulator, and $r(s_{t},a_{t})$ is the reward function. The goal is to learn a policy $\pi_{\theta}(a_{t}\mid s_{t})$ that maximizes the expected discounted return:
@@ 3.2 Problem Formulation
3) Action space: The action $a_{t}\in\mathbb{R}^{12}$ specifies the desired joint position targets for the robot’s 12 actuated joints. A proportional–derivative (PD) controller [tan2009computation] converts these position targets into torque commands applied to the actuators.
@@
3.3 SSD-Mamba2 Fusion Backbone
@@
Figure 2: Schematic of a SSD-Mamba2 Layer.
@@
3.4 Policy Optimization with PPO
PPO is a widely used model-free reinforcement learning algorithm for continuous control due to its stability and simplicity. Instead of performing large, potentially destabilizing policy updates, PPO restricts the update step with a clipping mechanism, ensuring stable and efficient training [schulman2017proximal]. The objective of PPO is defined as:
@@ 3.4 Policy Optimization with PPO
3.5 Safety Considerations
We consider safety during both training and execution, with the goal of improving empirical robustness.
During training, several mechanisms are employed to mitigate unsafe exploration. Episodes are terminated immediately upon falling or severe collisions, which prevents repeated exposure to hazardous states. The policy outputs bounded joint position targets that are executed through a low-level PD controller, thereby reducing abrupt or unstable actuation. In addition, an obstacle-density curriculum gradually increases task difficulty, enabling the agent to acquire basic locomotion skills before encountering complex or cluttered environments. Domain randomization and visual perturbations further introduce variations in terrain properties, physical parameters, and sensor characteristics, exposing the policy to a wide range of conditions and reducing overfitting.
At execution time, the learned policy operates at a high level by producing joint position targets rather than raw torque commands. This design makes it possible to combine the policy with conventional low-level control and safety mechanisms, such as joint limits, collision monitors, or emergency stop systems. Moreover, the bounded and input-gated state-space updates in the SSD-Mamba2 backbone suppress spurious sensory inputs, contributing to stable behavior under observation uncertainty.
4 Experiments
4.1 Environment Setup
All experiments are carried out on a laptop with an Intel Xeon(R) Platinum 8358P CPU (128 cores, 2.6 GHz base frequency) and a NVIDIA GeForce RTX 3090 GPU (24 GB, CUDA 12.8). The computing environment is based on Ubuntu 22.04, with physics simulation executed in PyBullet [coumans2021pybullet]. All models are developed in Python 3.8 using PyTorch 2.4.1.
@@ 4.1 Environment Setup
Figure 3: Simulated environments. Panels (a)–(c): (a) Thin Obstacles with Goals; (b) Thin Obstacle with Rugged Terrain; (c) Sphere Obstacles with Goals. Obstacle layouts are randomized at reset; all obstacle positions are static during an episode.
4.2 Baseline Methods